OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. The sum of two numbers is 170 and their difference is 26. Find the two numbers. Show how to solve this using any method Please help <3

OpenStudy (gw2011):

Here is the solution even though it is late (better late than never): x+y=170 x-y=26 Add these two equations to get: 2x=196 x=98 substitute this value into one of the equations to get: x+y=170 98+y=170 y=170-98 y=72 So the answer is: x=98 and y=72